Calculate the durations and volatilities of securities A, B, and C. Their cash flows are shown below. The interest rate is 8%. (Do not round intermediate calculations. Round "Duration" to 4 decimal places and "Volatility" to 2 decimal places.)

Period 1 Period 2 Period 3 Duration Volatility
  A 40 40 40 years
  B 20 20 120 years
  C 10 10 110 years
Answer #1

Duration:-

Security A

YEAR (1) CASH FLOW (2) PRESENT VALUE FACTOR@ 8% (3) PRESENT VALUE (PV) (4) PROPORTION OF TOTAL PV (5) DURATION = PROPORTION * TIME [(6) = (5) * (1)]
1 40 0.9259 37.036 0.3593 0.3593
2 40 0.8573 34.292 0.3327 0.6654
3 40 0.7938 31.752 0.3080 0.9240
TOTAL 103.08 1.0000 1.9487

DURATION = 1.9487 YEARS

Security B

YEAR (1) CASH FLOW (2) PRESENT VALUE FACTOR@ 8% (3) PRESENT VALUE (PV) (4) PROPORTION OF TOTAL PV (5) DURATION = PROPORTION * TIME [(6) = (5) * (1)]
1 20 0.9259 18.518 0.1414 0.1414
2 20 0.8573 17.146 0.1310 0.2620
3 120 0.7938 95.256 0.7276 2.1828
TOTAL 130.92 1.0000 2.5862

DURATION = 2.5862 YEARS

Security C

YEAR (1) CASH FLOW (2) PRESENT VALUE FACTOR@ 8% (3) PRESENT VALUE (PV) (4) PROPORTION OF TOTAL PV (5) DURATION = PROPORTION * TIME [(6) = (5) * (1)]
1 10 0.9259 9.259 0.0881 0.0881
2 10 0.8573 8.573 0.0815 0.1630
3 110 0.7938 87.318 0.8304 2.4912
TOTAL 105.15 1.0000 2.7423

DURATION = 2.7423 YEARS

VOLATILITY :-

VOLATILITY = DURATION / (1+YIELD)

YIELD = 8%(GIVEN)

SECURITY A = 1.9487/1.08 = 1.80

SECURITY B = 2.5862/1.08 = 2.39

SECURITY C = 2.7423/1.08 = 2.54